A product sells for $210 per unit, and its variable costs are 60% of sales. The fixed costs are $408,000. What is the break-even point in sales dollars? (Do not round intermediate calculations.)

Multiple Choice

  • $680,000.

  • $1,943.

  • $4,857.

  • $408,000.

  • $1,020,000.
Answer #1

Contribution margin =Sales-Variable cost

=(100-60)%=40%

Hence breakeven sales=Fixed cost/Contribution margin ratio

=(408,000/0.4)

=$1,020,000.